# A LLM-Based Ranking Method for the Evaluation of Automatic Counter-Narrative Generation

### v1

@@ 4.2 Feature Evaluation
For a comprehensive evaluation of CN quality, we conducted a manual assessment based on five criteria: Relatedness, Specificity, Richness, Coherence, Grammaticality and overall score Bengoetxea et al. (2024). Detailed evaluation guidelines can be found in Appendix A.2. The evaluation was conducted by 2 annotators on 90 HS-CN pairs, which included 10 HS instances from the CONAN corpus and the corresponding CNs generated by each of the 9 models. Each evaluated feature used a five-point scale for the answers, with 1 as the lowest score and 5 as the highest. The mean IAA across all the evaluated features is $0.69$. The feature-wise IAA is presented in Appendix B.
Model Version Type lrCONAN lrMT-CONAN
mistral v0.1 Base 1e-5 1e-4
mistral-instruct v0.2 Instruct 6e-6 3e-5
zephyr Beta Chat 6e-6 1e-4
llama-chat llama 2 Chat 2e-5 1e-3
@@ 5 Experimental Setup
This section outlines the specific setup employed for the generation of CNs. We consider two generation scenarios: ZS (see Section 5.1) and inference with fine-tuned models (see Section 5.2).
@@ 5.2 Fine-Tuning Details
Rank Human Score JudgeLM Score
@@
6 mistral 9.08 mistralzs 9.05
7 mistral-instruct 7.54 mistral 8.70
8 llama-chatzs 7.26 mistral-instruct 8.50
9 llama-chat 3.35 llama-chat 4.11
Table 4: Comparison of human and JudgeLM rankings, including the final scores obtained from the pairwise tournaments. ${zs}$ means that the model was used in a zero-shot setting.
@@ 6.1 Correlation of Automatic Metrics with Human Ratings
mistral-instruct 3.75 3.55 3.30 3.10 4.30 2.70
mistral 3.65 3.55 3.05 3.30 4.35 2.60
zephyr 4.40 4.75 3.60 3.20 4.35 2.30
llama-chat 3.40 3.10 2.95 3.30 4.10 2.20
@@ 7 Analysis
To confirm which of the models from Section 3.1 is the best for the task, we conducted a final feature-wise evaluation as explained in Section 4.2. The results are presented in Table 5. As seen there, the best-performing model is undoubtedly Zephyr, which considerably surpasses the gold standard.
Analyzing Table 4 in more detail, which presents the final rankings obtained from manual and automatic pairwise comparisons, we note that while manual evaluators were instructed to select a winning CN unless both were deemed ineffective, JudgeLM assigns ties when both responses are of high quality. This approach may result in a lower correlation between JudgeLM ratings and human preference.
@@ 8 Conclusions
CN generation needs a different evaluation framework and metrics than those used in previous work Chung et al. (2021); Tekiroğlu et al. (2022); Bengoetxea et al. (2024). This is due to the unique objectives, complexities, and impact of CNs, which require specialized criteria to assess their effectiveness and quality accurately. Thus, developing and implementing tailored evaluation metrics is crucial to advance the field and ensure the successful creation of impactful CNs.
Consistent with previous research observations, traditional metrics fall short in evaluating generation tasks that require creativity, including CN generation to combat HS. In this paper we present a LLM-based ranking method to provide an alternative automatic evaluation technique that offers promise, as it has shown increased correlation with human evaluation.
@@
In future work, we aim to extend this analysis to other languages such as Spanish, along with Basque, which is considered a low-resource language. We also intend to explore Retrieval Augmented Generation (RAG) to address the truthfulness issue, as we anticipate that this approach could substantially enhance the correlation between human evaluations and those conducted by Judge Models.

### v3

@@ 4.2 Feature Evaluation
For a comprehensive evaluation of CN quality, we provide a manual assessment based on five criteria: Relatedness, Specificity, Richness, Coherence, Grammaticality and overall score (Bengoetxea et al., 2024). Detailed evaluation guidelines can be found in Appendix A.2. The evaluation involved 2 annotators on 90 HS-CN pairs, which included 10 HS instances from the CONAN corpus and the corresponding CNs generated by each of the 9 models. Each evaluated feature used a five-point scale for the answers, with 1 as the lowest score and 5 as the highest. The mean IAA across all the evaluated features is $0.69$. The feature-wise IAA is presented in Appendix B.
@@ 5 Experimental Setup
This section outlines the specific setup employed for the generation of CNs. We consider two generation scenarios: ZS inference (see Section 5.1), and inference with fine-tuned models (see Section 5.2).
@@ 5.2 Fine-Tuning Details
Rank Human Score JudgeLM33B Score
@@
6 mistralft 9.08 mistralzs 9.05
7 mistral-instructft 7.54 mistralft 8.70
8 llama-chatzs 7.26 mistral-instructft 8.50
9 llama-chatft 3.35 llama-chatft 4.11
Table 4: Comparison of human and JudgeLM rankings, including the final scores obtained from the pairwise tournaments. ${\mathrm{zs}}$ means that the model was used in a zero-shot setting.
@@ 6.1 Correlation of Automatic Metrics with Human Ratings
mistral-instructft 3.75 3.55 3.30 3.10 4.30 2.70
mistralft 3.65 3.55 3.05 3.30 4.35 2.60
zephyrft 4.40 4.75 3.60 3.20 4.35 2.30
llama-chatft 3.40 3.10 2.95 3.30 4.10 2.20
@@ 7 Analysis
To confirm which of the models from Section 3.1 is the best for the task, we undertook a final feature-wise evaluation as explained in Section 4.2. The results are presented in Table 5. As seen there, the best-performing model is undoubtedly Zephyr, which considerably surpasses the gold standard.
Analyzing Table 4 in more detail, which presents the final rankings obtained from manual and automatic pairwise comparisons, we note that while manual evaluators were instructed to select a winning CN unless both were deemed ineffective, JudgeLM assigns ties when both responses are of high quality. This approach may result in a lower correlation between JudgeLM ratings and human preference. Upon a more nuanced analysis, we also observed that JudgeLM demonstrates a distinct preference for factual CNs, particularly those offering detailed and specific information about research and related topics. However, the model lacks the capability to reliably detect hallucinations, meaning it may favor responses that appear factual without verifying the truthfulness or accuracy of the provided details.
@@ 8 Conclusions
CN generation needs a different evaluation framework and metrics than those used in previous work (Chung et al., 2021; Tekiroğlu et al., 2022; Bengoetxea et al., 2024). This is due to the unique objectives, complexities, and impact of CNs, which require specialized criteria to assess their effectiveness and quality accurately. Thus, developing and implementing tailored evaluation metrics is crucial to advance the field and ensure the successful creation of impactful CNs.
Consistent with previous research observations, traditional metrics fall short in evaluating generation tasks that require creativity, including CN generation to combat HS. In this paper we present a promising LLM-based ranking method to provide an alternative automatic evaluation technique which exhibits higher correlation with human evaluation.
@@
In future work, we aim to extend this analysis to other languages such as Spanish, along with Basque, which is considered a low-resource language. Finally, we plan to explore Retrieval Augmented Generation (RAG) to address the truthfulness issue, as we anticipate that this approach could substantially enhance the correlation between human evaluations and those of Judge Models.